# Personalization of Large Language Models: A Survey

### v1

@@ 3.3.4 Pair-Wise Human Preferences
Figure 3: Space of Personalized Generations. We characterize the space of generations for query $x$, including the space of all possible generations, the space of all high quality generations, and finally, the space of user-specific personalized generations for user $i$. Intuitively, given two users $i$ and $j$, the space of high-quality personalized generations for each user may be completely disjoint.
@@ 3.4 Space of Personalized Generations
In this section, we briefly formalize and analyze the problem of personalized LLMs and their solution spaces. This serves two purposes: providing intuition on the difficulty of the problem and characterizing the properties and unique advantages that relate to other well-studied problems.
Let us first establish the formalization of the personalized LLM problem. Consider a generic input example $x\in\mathcal{X}$. We denote the generative model by $g:\mathcal{Z}\times\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{Z}$ represents the latent space and $\mathcal{Y}$ represents the space of all possible generations.
Given $x\in\mathcal{X}$, the space of all possible generations is defined as:
@@
    $\mathcal{S}_{i}(x)=\{y\in\mathcal{S}_{h}(x):f(P_{u_{i}},y)\leq\epsilon,P(y|x)\geq\delta\}\}$
@@
Note that the space of user-specific generations $\mathcal{S}_{i}(x)$ is significantly smaller and more targeted compared to the space of all possible generations $\mathcal{S}(x)$. An intuitive overview of the space of personalized generations for a specific user can be found in Figure 3.
@@ 3.5 Personalization Criterion Taxonomy
When evaluating the personalization of generated text in LLMs, it is essential to consider several critical aspects to ensure the content is effectively tailored to individual users. These aspects constitute a comprehensive taxonomy of personalization criteria, encompassing various dimensions of personalized content generation.
@@ 3.6.3 Taxonomy of Evaluation Methodologies for Personalized LLMs
Definition 12 (Intrinsic Evaluation).
$\mathbb{E}_{i}(\hat{y},Y)=\psi(\hat{y},Y)$
Definition 13 (Extrinsic Evaluation).
$\mathbb{E}_{e}(\hat{r},r)=\psi_{a}(\hat{r},r)$
@@ 3.6.4 Taxonomy of Datasets for Personalized LLMs
$\bm{\mathrm{v}}$ task-specific information
@@
$\mathcal{H_{\text{sys}}}$ the system prompt to input LLMs
$\mathcal{H_{\text{usr}}}$ the user prompt to input LLMs
@@
$Emb(\cdot)$ word or sentence encoder, which can be a part of $\mathcal{M}$
@@
$\bm{\mathrm{z}}$ embedding generated from user $i$’s information
@@ 4 Personalization Granularity of LLMs
  • User-level Personalization (Sec. 4.1): This level focuses on the unique preferences and behaviors of a single user. Personalization at this level utilizes detailed information about the user, including their historical interactions, preferences, and behaviors, often identified through a user ID (Li et al., 2024e). Formally, let $U$ represent the set of users, and $P_{u}=\{p_{u}^{1},p_{u}^{2},...,p_{u}^{n}\}$ denote the set of personalized preferences for user $u\in U$. The objective function of the downstream task is $\mathcal{L}_{task}$. The objective of personalization on this level is to minimize this function:
@@ 4.1 User-level Personalization
In this section, we discuss user-level personalization, which focuses on data at the individual level. As depicted in Figure 6(a), this type of personalization focuses on optimizing preferences for each user uniquely identified by a user ID. For instance, in the MovieLens-1M recommendation dataset (Harper & Konstan, 2015), each user has demographic information such as UserID, Gender, Age, Occupation, and Zip-code, alongside corresponding movie interactions (MovieID, Rating, Timestamp). The goal is to recommend new movies based on each user’s profile and viewing history. The advantage of this level of personalization is that it offers the most fine-grained approach, minimizing noise from other users. This is particularly beneficial in domains such as online shopping, job recommendations (Wu et al., 2024), and healthcare (Abbasian et al., 2023; 2024; Zhang et al., 2024a; Jin et al., 2024b), where individual user behavior can vary significantly, and such detailed, individualized personalization is crucial. One of the main challenges of this level of personalization is the “cold-start problem” which refers to users with minimal interaction history, often termed “lurkers” in recommendation systems (Sun et al., 2024). However, many studies (Salemi et al., 2023; Rajput et al., 2023; Xi et al., 2023) choose to remove such data during the preprocessing stages. This exclusion potentially undermines the robustness of the systems by disregarding the subtleties and potential insights offered by these underrepresented user interactions.
@@ 4.2 Persona-level Personalization
In this section, we formalize and discuss persona-level personalization, where the input comprises the preferences of users categorized by group or persona. As illustrated in Figure 6(b), this approach targets optimizing the preferences of a user group sharing common characteristics. A natural language description encapsulating these shared traits represents the entire group within prompts or other relevant components. For example, Jang et al. (2023) design three distinct perspectives of preferences: expertise, informativeness, and style, with each dimension featuring two conflicting personas or preferences. For instance, in the expertise dimension, one persona prefers content that is easily understandable by an elementary school student, while the other prefers content that is comparable only to a PhD student in the specific field. From this example, we can observe that, compared to localized user-specific personalization (Sec. 4.1), each persona represents a broader portrait of a group of users, focusing on more general features rather than detailed user-specific information. The advantage of persona-level personalization lies in its effectiveness in scenarios where shared characteristics are prominent and crucial for downstream tasks, while user-specific attributes are less significant. Additionally, once these characteristics are extracted, this data format is easier to process, either by including it directly in the prompt or utilizing it through RLHF, compared to lengthy user-specific profiles. However, the challenge is that extracting such representative characteristics using natural language description can be difficult in practice. Most current works still rely on human’s domain knowledge to achieve this.
4.3 Global Preference Alignment Personalization
@@
In many applications, only global user preference data may be available, representing the preferences of the entire population rather than those of individual users. While this falls outside the primary scope of personalization in this survey, we include a discussion of it for completeness. These preferences typically encompass human values expected to be accepted by the general public, such as social norms, factual correctness, and instruction following (Taylor et al., 2016; Gabriel, 2020; Liu, 2024). The common format of such data includes a given instruction, multiple options, and a label annotated by human annotators indicating which option is preferable (Ethayarajh et al., 2022; Stiennon et al., 2020a; Nakano et al., 2021; Bai et al., 2022; Ganguli et al., 2022). These datasets are typically used through RLHF to align LLMs. The advantage of global preference alignment is its potential to enhance LLMs in terms of safety (Gehman et al., 2020; Ge et al., 2023; Anwar et al., 2024; Ji et al., 2024a), social norms (Ryan et al., 2024), and ethical issues (Liu et al., 2021; Rao et al., 2023), ensuring they align with human values. However, the disadvantage is that it may introduce noise, as individual preferences can vary and may not always represent the general public accurately. Moreover, this level of alignment does not capture fine-grained, user-specific personalization.
@@ 5 Taxonomy of Personalization Techniques for LLMs
  • Personalization via Prompting (Sec 5.2): This category of methods incorporates user information as the context within the prompts for LLMs. By providing this contextual information, LLMs can either directly perform downstream personalization tasks through text generation or act as intermediate modules to extract more relevant information, thereby enhancing the system’s performance.
@@ 5.1 Personalization via Retrieval-Augmented Generation
Retrieval-augmented generation (RAG) enhances LLM performance by retrieving relevant document segments from an external knowledge base using semantic similarity calculations (Gao et al., 2023). This approach is widely employed in information retrieval and recommendation systems (Zhao et al., 2024b; Rajput et al., 2023; Di Palma, 2023; Wang et al., 2024f). While RAG can reduce hallucinations by grounding the generation process in retrieved factual content (Shuster et al., 2021; Li et al., 2024c), these retrieval modules can also be used to retrieve personalized information, enabling the generation of customized, tailored outputs.
@@
Retrieval-augmented Generation (RAG) is a process in which a language model $\mathcal{M}$ leverages a retrieval model $\mathcal{R}$ to enhance its generation capabilities. Given an input $X_{i}$ from user $i$, the retrieval model $\mathcal{R}$ identifies $k$ relevant external data points or documents from a dataset $\mathcal{D}$. These retrieved data points are then incorporated into the input text to form a transformed input $\bar{x}$, which is used by the language model $\mathcal{M}$ to generate a grounded output $\hat{y}_{i}$.
@@
$\bar{x}=\phi_{p}(X_{i},\mathcal{R}(\phi_{q}(X_{i}),\mathcal{D},k))$
where $\phi_{q}$ is a query construction function used by the retrieval model $\mathcal{R}$ to find relevant documents, and $\phi_{p}$ is a prompt construction function that integrates the retrieved information into the original input $X_{i}$. The output $\hat{y}_{i}$ represents the generated text based on both the original input and the retrieved information.
For personalization tasks, large user profiles often act as the external knowledge base because they cannot be fully incorporated into the prompt due to LLMs’ context limit. Li et al. (2023b) propose a multi-stage multi-task framework containing retrieval, ranking, summarization, synthesis, and generation to conduct personalized document generation. Salemi et al. (2023) use both sparse and dense retrievers to retrieve relevant user information to perform personalized classification and generation tasks. Mysore et al. (2023a) personalizes LLM writing assistants by using a generation-calibrated retriever, which selects historical user-authored documents to enhance prompt augmentation for better personalization. They include a training data selection method and a scale-calibrating KL-divergence objective to optimize the retriever’s effectiveness.
@@ 5.1.1 Sparse Retrieval
BM25: BM25 is a more advanced sparse retrieval method that extends the TF-IDF model by incorporating document length normalization and saturation controls for term frequency. The BM25 score for a document $D$ with respect to a query $q_{i}$ is calculated as:
@@
Because of their generalizability, effectiveness, and simplicity, sparse retrievers often serve as baselines in retrieval-based personalization methods (Salemi et al., 2023; Li et al., 2023b; Richardson et al., 2023). For instance, Salemi et al. (2023) uses BM25 as one of the retrievers to fetch relevant user information, which is then incorporated into prompts for LLMs when evaluating on the LaMP dataset (Salemi et al., 2023). Richardson et al. (2023) enhance personalization in LLMs by integrating BM25 retrieval with user data summaries, achieving improved performance on LaMP tasks while reducing the volume of retrieved data by 75%. In another work, Li et al. (2023b) proposes a multistage, multitask framework inspired by writing education to enhance personalized text generation in LLMs. In the initial retrieval stage, BM25 is used to retrieve relevant past user documents, which are subsequently ranked and summarized to generate personalized text.
@@ 5.1.2 Dense Retrieval
Dense retrievers leverage deep neural networks to generate continuous representations for queries and documents, enabling retrieval in a dense embedding space via similarity-based search (Johnson et al., 2021). Some works (Sun et al., 2024) utilize pre-trained LLM encoders, such as OpenAI’s text-embedding-ada series and Sentence-BERT (Reimers & Gurevych, 2019), without fine-tuning their parameters. Other approaches focus on training retrieval-oriented embeddings. For example, Dense Passage Retrieval (DPR) (Karpukhin et al., 2020) employs dense embeddings in a dual-encoder framework, using BM25 hard negatives and in-batch negatives to efficiently retrieve relevant passages for open-domain question answering. Contriever (Izacard et al., 2021) is an unsupervised dense retriever trained using contrastive learning, where two random spans from a document are cropped independently to form positive pairs for training. In the context of personalization, several works have proposed specialized training data construction (Mysore et al., 2023a; b) and training strategies (Mysore et al., 2023a; Salemi et al., 2023; 2024) to enhance dense retrievers’ ability to retrieve more relevant user information, improving personalization for downstream tasks using LLMs. Salemi et al. (2023) use Fusion-in-Decoder (Izacard & Grave, 2021) on encoder-decoder models such as T5 Raffel et al. (2020) which retrieves and concatenate multiple relevant documents’ encoded embedding before decoding. (Mysore et al., 2023a) train a pre-trained MPNET model (Song et al., 2020) with a scale-calibrating KL-divergence objective which shows superior performance on personalized open-ended long-form text generation tasks. Salemi et al. (2024) train the Contriever model for personalizing LLMs using LLMs’ feedback with policy gradient optimization and knowledge distillation. Zeng et al. (2023) introduces UIA, a flexible dense retrieval framework that incorporates personalized attentive networks to enhance various information access such as keyword search, query by example, and complementary item recommendation. Other dense retrievers such as Sentence-T5 (Ni et al., 2021) and Generalizable T5-based dense Retrievers (GTR) (Ni et al., 2022) are also frequently used for downstream personalization tasks. Generally, while dense retrievers require training on downstream tasks, making them more costly and time-inefficient (Richardson et al., 2023), they tend to achieve superior performance compared to sparse retrievers in downstream personalization tasks. However, constructing effective training data, designing suitable loss functions, and incorporating LLMs into the training process to optimize retrievers for improved downstream personalization tasks remain open challenges.

### v3

@@ 3.3.4 Pair-Wise Human Preferences
Figure 3: Space of Personalized Generations. We characterize the space of generations for query $x$, including the space of all possible generations $\mathcal{S}(x)$, the space of all high quality generations $\mathcal{S}_{h}(x)$, and finally, the space of user-specific high-quality personalized generations $\mathcal{S}_{i}(x)$ for user $i$. Intuitively, given two users $i$ and $j$, the space of high-quality personalized generations for each user may be completely disjoint.
@@ 3.4 Space of Personalized Generations
In this section, we briefly formalize and analyze the problem of personalized LLMs and their solution spaces and provide an intuitive overview in Figure 3. This serves two purposes: providing intuition on the difficulty of the problem and characterizing the properties and unique advantages that relate to other well-studied problems.
Let us first establish the formalization of the personalized LLM problem. Consider a generic input example $x\in\mathcal{X}$. We denote the generative model by $g:\mathcal{Z}\times\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{Z}$ represents the latent space and $\mathcal{Y}$ represents the space of all possible generations. Given input $x\in\mathcal{X}$, the space of all possible generations is
@@
  • The space of all possible generations for a given input $x$ is $\mathcal{S}(x)$.
@@
    $\mathcal{S}_{i}(x)=\{y\in\mathcal{S}_{h}(x):f(P_{u_{i}},y)\leq\epsilon,P(y|x)\geq\delta\}$
@@
An intuitive overview of the space of personalized generations for a specific user can be found in Figure 3. Note that the space of user-specific generations $\mathcal{S}_{i}(x)$ is significantly smaller and more targeted compared to the space of all possible generations $\mathcal{S}(x)$. One example is that there may be many correct responses to a specific question, however, only a very small subset of answers may capture the important details needed for the answer to be useful to the specific user. In particular, the user may need additional steps to carry out a specific task, or they may know certain terminology and language that would enable better understanding for the user. Overall, the key takeaways are that the user personalization tasks are extremely challenging, though will only become increasingly important in the future. In particular, there is only a very tiny space of responses that are useful to a specific user, and generating such a response is only becoming more and more important. This highlights the need to not only develop better techniques to generate such user-specific responses, but also better data and evaluation metrics to quantify it.
@@ 3.5 Personalization Criterion Taxonomy
When evaluating the personalization of generated text in LLMs, it is essential to consider several critical aspects to ensure the content is effectively tailored to individual users. These aspects constitute a taxonomy of personalization criteria, encompassing various dimensions of personalized content generation.
@@ 3.6.4 Taxonomy of Datasets for Personalized LLMs
$\bm{\mathrm{v}}$ task-specific feature vector
@@
$\mathcal{H_{sys}}$ The system prompt to input LLMs
$\mathcal{H_{usr}}$ The user prompt to input LLMs
@@
$E(\cdot)$ word or sentence encoder, which can be a part of $\mathcal{M}$
@@
$\bm{\mathrm{z}}$ embedding of generated text $\hat{Y}_{i}$ for user $i$
$\bm{\mathrm{r}}$ the output of a personalized system for a downstream task
@@ 4 Personalization Granularity of LLMs
  • User-level Personalization (Sec. 4.1): This level focuses on the unique preferences and behaviors of a single user. Personalization at this level utilizes detailed information about the user, including their historical interactions, preferences, and behaviors, often identified through a user ID (Li et al., 2024g). Formally, let $U$ represent the set of users, and $P_{u}=\{p_{u}^{1},p_{u}^{2},...,p_{u}^{n}\}$ denote the set of personalized preferences for user $u\in U$. The objective function of the downstream task is $\mathcal{L}_{task}$. The objective of personalization on this level is to minimize this function:
@@ 4.1 User-level Personalization
In this section, we discuss user-level personalization, which focuses on data at the individual level (Zollo et al., 2024). As depicted in Figure 6(a), this type of personalization focuses on optimizing preferences for each user uniquely identified by a user ID. For instance, in the MovieLens-1M recommendation dataset (Harper & Konstan, 2015), each user has demographic information such as UserID, Gender, Age, Occupation, and Zip-code, alongside corresponding movie interactions (MovieID, Rating, Timestamp). The goal is to recommend new movies based on each user’s profile and viewing history. The advantage of this level of personalization is that it offers the most fine-grained approach, minimizing noise from other users. This is particularly beneficial in domains such as online shopping, job recommendations (Wu et al., 2024b), and healthcare (Abbasian et al., 2023; 2024; Zhang et al., 2024a; Jin et al., 2024b), where individual user behavior can vary significantly, and such detailed, individualized personalization is crucial. One of the main challenges of this level of personalization is the “cold-start problem” which refers to users with minimal interaction history, often termed “lurkers” in recommendation systems (Sun et al., 2024). However, many studies (Salemi et al., 2023; Rajput et al., 2023; Xi et al., 2023) choose to remove such data during the preprocessing stages. This exclusion potentially undermines the robustness of the systems by disregarding the subtleties and potential insights offered by these underrepresented user interactions.
@@ 4.2 Persona-level Personalization
In this section, we discuss persona-level personalization, where the input comprises the preferences of users categorized by group or persona. As illustrated in Figure 6(b), this approach targets optimizing the preferences of a user group sharing common characteristics. A natural language description encapsulating these shared traits represents the entire group within prompts or other relevant components. For example, Jang et al. (2023) design three distinct perspectives of preferences: expertise, informativeness, and style, with each dimension featuring two conflicting personas or preferences. For instance, in the expertise dimension, one persona prefers content that is easily understandable by an elementary school student, while the other prefers content that is comparable only to a PhD student in the specific field. From this example, we can observe that, compared to localized user-specific personalization (Sec. 4.1), each persona represents a broader portrait of a group of users, focusing on more general features rather than detailed user-specific information. The advantage of persona-level personalization lies in its effectiveness in scenarios where shared characteristics are prominent and crucial for downstream tasks, while user-specific attributes are less significant. Additionally, once these characteristics are extracted, this data format is easier to process, either by including it directly in the prompt or utilizing it through RLHF, compared to lengthy user-specific profiles. However, extracting representative characteristics through natural language descriptions can be challenging in practice, often requiring substantial reliance on human domain knowledge.
4.3 Global Preference Personalization
@@
In many applications, only global user preference data may be available, representing the preferences of the entire population rather than those of individual users. While this falls outside the primary scope of personalization in this survey, we include a discussion of it for completeness. These preferences typically encompass human values expected to be accepted by the general public, such as social norms, factual correctness, and instruction following (Taylor et al., 2016; Gabriel, 2020; Liu, 2024). The common format of such data includes a given instruction, multiple options, and a label annotated by human annotators indicating which option is preferable (Ethayarajh et al., 2022; Stiennon et al., 2020a; Nakano et al., 2021; Bai et al., 2022; Ganguli et al., 2022). These datasets are typically used through RLHF to align LLMs. The advantage of global preference alignment is its potential to enhance LLMs in terms of safety (Gehman et al., 2020; Ge et al., 2023; Anwar et al., 2024; Ji et al., 2024a), social norms (Ryan et al., 2024), and ethical issues (Liu et al., 2021; Rao et al., 2023), ensuring they align with human values. However, the disadvantage is that it may introduce noise, as individual preferences can vary and may not always represent the general public accurately. Moreover, this level of alignment does not capture fine-grained personalization.
@@ 5 Taxonomy of Personalization Techniques for LLMs
  • Personalization via Prompting (Sec 5.2): This category of methods incorporates user information as the context within the prompts for LLMs. By providing this contextual information, LLMs can either directly perform downstream personalization tasks through text generation or act as intermediate modules to extract more relevant information, thereby enhancing the system’s performance on downstream tasks.
@@
Figure 7: A Case Study on Personalized Movie Review Generation. This figure illustrates how different personalization techniques for LLMs enhance personalized movie review generation by leveraging user profiles, retrieval modules, and fine-tuning methods to align outputs with individual preferences and writing styles.
@@ 5.1 Personalization via Retrieval-Augmented Generation
Retrieval-augmented generation (RAG) enhances LLM performance by retrieving relevant document segments from an external knowledge base using semantic similarity calculations (Gao et al., 2023). This approach is widely employed in information retrieval and recommendation systems (Zhao et al., 2024b; Rajput et al., 2023; Di Palma, 2023; Wang et al., 2024g). While RAG can reduce hallucinations by grounding the generation process in retrieved factual content (Shuster et al., 2021; Li et al., 2024d), these retrieval modules can also be used to retrieve personalized information, enabling the generation of customized, tailored outputs.
@@
Retrieval-augmented Generation (RAG) is a process in which a language model $\mathcal{M}$ leverages a retrieval model $\mathcal{R}$ to enhance its generation capabilities. Given an input $X_{u}$ from user $u$, the retrieval model $\mathcal{R}$ identifies $k$ relevant external data points or documents from a dataset $\mathcal{D}$. These retrieved data points are then incorporated into the input text to form a transformed input $\bar{x}$, which is used by the language model $\mathcal{M}$ to generate a grounded output $\hat{y}_{u}$.
@@
$\bar{x}=\phi_{p}(X_{u},\mathcal{R}(\phi_{q}(X_{u}),\mathcal{D},k))$
$\hat{y}_{i}=M(\bar{x})$
where $\phi_{q}$ is a query construction function used by the retrieval model $\mathcal{R}$ to find relevant documents, and $\phi_{p}$ is a prompt construction function that integrates the retrieved information into the original input $X_{u}$. The output $\hat{y}_{u}$ represents the generated text based on both the original input and the retrieved information.
@@ 5.1.1 Sparse Retrieval
BM25: This is a more advanced sparse retrieval method that extends the TF-IDF model by incorporating document length normalization and saturation controls for term frequency. The BM25 score for a document $D$ with respect to a query $q_{i}$ is calculated as:
@@
Because of their generalizability, effectiveness, and simplicity, sparse retrievers often serve as baselines in retrieval-based personalization methods (Salemi et al., 2023; Li et al., 2023b; Richardson et al., 2023). For instance, Salemi et al. (2023) use BM25 as one of the retrievers to fetch relevant user information, which is then incorporated into prompts for LLMs when evaluating on the LaMP dataset (Salemi et al., 2023). Richardson et al. (2023) enhance personalization in LLMs by integrating BM25 retrieval with user data summaries, achieving improved performance on LaMP tasks while reducing the volume of retrieved data by 75%. In another work, Li et al. (2023b) propose a multistage, multitask framework inspired by writing education to enhance personalized text generation in LLMs. In the initial retrieval stage, BM25 is used to retrieve relevant past user documents, which are subsequently ranked and summarized to generate personalized text.
@@ 5.1.2 Dense Retrieval
Dense retrievers leverage deep neural networks to generate continuous representations for queries and documents, enabling retrieval in a dense embedding space via similarity-based search (Johnson et al., 2021). Some works (Sun et al., 2024) employ pre-trained LLM encoders, such as OpenAI’s text-embedding-ada series and Sentence-BERT (Reimers & Gurevych, 2019), without fine-tuning their parameters. Other approaches focus on training retrieval-oriented embeddings. For example, Dense Passage Retrieval (DPR) (Karpukhin et al., 2020) employs dense embeddings in a dual-encoder framework, using BM25 hard negatives and in-batch negatives to efficiently retrieve relevant passages for open-domain question answering. Contriever (Izacard et al., 2021) is an unsupervised dense retriever trained using contrastive learning, where two random spans from a document are cropped independently to form positive pairs for training. In the context of personalization, several works have proposed specialized training data construction (Mysore et al., 2023a; b) and training strategies (Salemi et al., 2023; 2024) to enhance dense retrievers’ ability to retrieve more relevant user information, improving personalization for downstream tasks using LLMs. Salemi et al. (2023) use Fusion-in-Decoder (Izacard & Grave, 2021) on encoder-decoder models such as T5 (Raffel et al., 2020) which retrieves and concatenate multiple relevant documents’ encoded embedding before decoding. Mysore et al. (2023a) train a pre-trained MPNET model (Song et al., 2020) with a scale-calibrating KL-divergence objective which shows superior performance on personalized open-ended long-form text generation tasks. Salemi et al. (2024) train the Contriever model for personalizing LLMs using LLMs’ feedback with policy gradient optimization and knowledge distillation. Zeng et al. (2023) introduce UIA, a flexible dense retrieval framework that incorporates personalized attentive networks to enhance various information access such as keyword search, query by example, and complementary item recommendation. Other dense retrievers such as Sentence-T5 (Ni et al., 2021) and Generalizable T5-based dense Retrievers (GTR) (Ni et al., 2022) are also frequently used for downstream personalization tasks. Generally, while dense retrievers require training on downstream tasks, making them more costly and time-inefficient (Richardson et al., 2023), they tend to achieve superior performance compared to sparse retrievers in downstream personalization tasks. However, constructing effective training data, designing suitable loss functions, and incorporating LLMs into the training process to optimize retrievers for improved downstream personalization tasks remain open challenges.